A bank receives a new demand deposit $10,000 and the legal reserve requirement is 20%.

Calculate:

i) the amount the bank can lend out

ii) the amount of money the banking system can potentially create
Answer #1

Ans

1 10000-20% of 10000=8000

2 Here money multiplier=1/reserve ratio=1/20%=5

So it can create=10000(5)=50000